Project

General

Profile

NDOS Commissioning Meeting 2011/05/09

Attendance: Jon, Peter, Kurt, Andrew, Minerba, Tricia, Susan, Tom, Sue, Jarek, Rick, Greg, Leon, Athans, Denis

Agenda:

  1. NDOS Status and Installation Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. AOB (All)

NDOS Status and Channel Recovery Plans

  • Denis: just oil filling; should be done this week.

DAQ/Electronics Update & Plans

  • Peter: need to push ahead with updating packages today
  • Jon: stress tests? Andrew: we've completed 1/4 so far. First test increased trigger rate; DAQ fell over around 140-160 Hz. Seems to be an issue with DDS. Tests will continue this week.
  • Leon: need to put issues into Redmine
  • Sue: would like to see DCMs re-enabled for logging timing buffer commands. Can be done after we roll forward. Will require monitoring.
  • Peter: firmware version of DCMs needs to be updated. Sue: we're not quite ready for that... maybe later this week after some tests are done.
  • Kurt: DAQAppMgr and DAQOperationsTools also to be updated later this week.

DCS Update & Plans

  • Athans: there were no tests this weekend because novatest01 crashed.
  • Sue: doesn't think there is a correlation between novatest01 crash and DCS tests. Tests have resumed as of this morning. It looks like there is a correlation between FEB readout and the DCM crashing during these tests.
  • Jon: there needs to be a meeting between DAQ and DCS groups soon to discuss integration between the two systems. Athans will coordinate a separate meeting.

AOB (All)

  • Lost ~24 hours of beam data because shifters were not paying attention to the trigger scalars. novatest01 had crashed, and this machine is currently a part of the beam triggers route.
  • Denis: we have not run pedestal runs.